OpenStudy (anonymous):

solve v^2+3v-54=0 6,-9??

OpenStudy (jamesj):

Again, you can verify by substituting. They look right to me.

OpenStudy (jamesj):

e.g., 6^2 + 3*6 - 54 = 36 + 18 - 54 = 0, so yes, 6 is a solution.

OpenStudy (anonymous):

yea you're right